-(3-10)-(6-3)
My answer is not coming out correctly.
Please see my version-
-(3+(-10))-3=
-(-7)-3=
-(-7)+(-3)= -13
The textbook gives the answer as 4. Where am I going wrong?
Thank you.

Answer by ptaylor(2198)

-(3-10)-(6-3)
My answer is not coming out correctly.
Please see my version-
-(3+(-10))-3=
-(-7)-3=
-(-7)+(-3)= -13
The textbook gives the answer as 4. Where am I going wrong?
Thank you.

YOU ARE CORRECT IN EVERY STEP EXCEPT THE LAST ONE:
-(-7)=+7 and +(-3) is -3

Hope this helps---ptaylor